# Thermodynamic inference in molecular motors: a Martingale approach

### v1

@@ I Introduction
One of the most important molecular motors in which we focus in this work is the ATP synthase complex [1], responsible for the synthesis of ATP, the energy currency of life. It appears in practically all living organisms, from prokaryotic cells to eukaryotic cells. The complex is composed of two main parts: the F0 domain, embedded in the cellular membrane, and the F1 domain, also called F1-ATPase. The latter is a rotary motor capable of using the energy released by translocation of protons across the F0 domain to synthesize or hydrolize ATP molecules [16, 17]. The F1-ATPase motor has been widely studied since the end of the 90s, both experimentally [18, 19] and theoretically [20]. The thermodynamic operation of the motor has been accurately explored in the last decade, assessing some of the main features of the motor and suggesting a nearly 100% thermodynamic efficiency in its transduction of chemical energy to mechanical work [21, 22]. In this context, the so-called fluctuation theorem (FT) [23] has been combined with single-molecule experiments to provide improved estimations of the rotatory torque in the motor [24], which constitutes one of the first applications of universal relations in stochastic thermodynamics for inference purposes [25, 26]. This inference concept [27] has also been applied and extended to many other active biological processes, and to other universal non-equilibrium relations, such as the relation between dissipation and irreversibility [28, 29, 30, 31, 32] , the thermodynamic uncertainty relation [33, 34, 35, 36], the thermodynamic speed limits [37, 38] or the infimum law of martingale theory [39, 40, 41].
@@ II Inference strategies using stopping times
In the following, we review relevant theoretical results on martingale theory for stochastic thermodynamics and introduce the two main inference strategies we propose in a generic configuration for discrete-state molecular motors performing cycles in its state space.
@@ II.1 IFT at stopping times
$e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta(\mathcal{T})}\right>_{\mathcal{T}}=1,$ (1)
@@
$e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta(\mathcal{T})}\right>_{\mathcal{T}}=1-\Gamma,$ (3)
@@ II.2 Counting cycles from observable transitions
Figure 1: Illustration of the stopping protocol introduced in the main text with sample trajectories from the F1-ATPase rotatory motor (see Sec. III). The current in the signalling transition $N(t)$ is monitored over time after the first transition is observed (after a time $t_{\rm onset}$), where we place the origin of times for the trajectories. Trajectories are stopped either when the current reaches the threshold value $N_{+}+1$ (blue trajectory) performing $N_{+}$ cycles in the forward direction, or $-N_{-}$ (orange trajectory) corresponding to $N_{-}$ cycles in the backward direction after the starting point at $t_{\rm onset}$.
The stopping condition we propose consists on a double threshold first-passage time for the current $N(t)$, as illustrated in Fig. 1: the dynamics stops when the net number of signaling transitions performed by the motor reaches either the upper threshold $N_{+}+1$ or the lower threshold $-N_{-}$, with positive numbers $N_{+}\neq N_{-}$ in general. We settle the starting point of the trajectories $X_{[0,\tau]}$ after the first signalling transition in the forward direction is detected, where we place the origin of time ($t_{\rm onset}$ in Fig. 1). We then let evolve the motor until the net number of jumps in the signalling transition reaches either the threshold $N_{+}+1$ or $-N_{-}$, which corresponds to having trajectories with $N_{+}$ net transitions in the forward direction (i.e. from the starting point) or $N_{-}+1$ transitions in the backward direction, respectively. Moreover, here we will consider a limit time $\tau$ large enough to guarantee that the stochastic dynamics always verify the stopping condition before reaching the limit time, i.e. $\mathcal{T}<\tau$. This condition is ensured since the functional $e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta_{\tau}(\mathcal{T})}$ remains bounded in the limit $\tau\rightarrow\infty$ [52], as guaranteed by the two-threshold first-crossing condition considered here.
@@
$e^{-S_{\mathrm{tot}}(\mathcal{T})}\right>_{\mathcal{T}}=P_{-}e^{-S_{\mathrm{tot}}^{-}}+P_{+}e^{-S_{\mathrm{tot}}^{+}}=1-\Gamma,$ (6)
@@
[…]
@@
On the other hand, when the boundary $-N_{-}$ is reached, the entropy production becomes:
@@
where $\Delta E_{x_{2}x_{1}}=E_{x_{1}}-E_{x_{2}}$ is the energy difference between mesostates $x_{1}$ and $x_{2}$, $\Delta N_{x_{2}x_{1}}^{\alpha}$ is the number of particles $\alpha$ exchanged with the reservoir in the transition, and $W_{x_{2}x_{1}}^{\rm ext}$ is a contribution to the mechanical work performed on the system by an external force pointing in the $x_{2}\rightarrow x_{1}$ direction.
@@
$e^{-S_{\mathrm{tot}}(\mathcal{T})}\right>_{\mathcal{T}}=$ $(1-P_{+})\pi_{x_{1}}e^{\beta\left[q_{x_{2}x_{1}}-N_{-}Q_{\mathrm{cycle}}\right]}$
@@
Remarkably, if the experimenter has access to the steady distribution values $\pi_{x_{1}}$ and $\pi_{x_{2}}$ (approximating $\Gamma\simeq 1-\pi_{x_{2}}$) and $q_{x_{1}x_{2}}$ is assumed to be known  [^1] In particular, when all transitions are observable, the local detail balance of each pair of transitions may be used to asses the energy (or free energy) changes between mesostates [27]. Moreover, the stationary distribution of the states could be obtained from the statistics of transitions, the heat per cycle $Q_{\mathrm{cycle}}$ can be inferred from Eq. (II.2), which may be used to obtain two main quantities: First, if the external forces acting on the system are known, one can infer the chemical work performed by the motor, and hence the corresponding differences in chemical potentials. Second, if all the chemical potentials are already known (as we will mainly consider in the following), one could use this information to measure external forces in the nanoscale.
@@ II.3 Inference from IFT with approximations
Nevertheless, in many situations of experimental interest only a few transitions among system states are observable. In that case, neither the stationary distribution of the states, $\pi_{x}$, nor the term $\Gamma$ would be accessible. To overcome this drawback, we propose a slightly different counting strategy and allow for some approximations. In this case, we assume to start the process directly from the stationary distribution $\pi_{x}$ in order to avoid absolute irreversibility. Therefore, now we start to count instances of the signalling transition without knowing exactly the initial state $x(0)$ (i.e. we do not wait any more for the first transition in the forward direction to happen to start counting). We then count the number of jumps in the signalling transition until the net flux $N(t)$ reaches one of the two thresholds $N_{+}$ or $-N_{-}$. We remark that in this scenario, both $\delta(\mathcal{T})$ and $\Gamma$ become zero. Besides, we approximate the entropy production at the two thresholds by:
@@ II.4 Bounds to maximal excursions of entropy production
In Ref. [45], universal families of bounds for the cumulative distribution of entropy production maxima and minima (and thus for the survival probability) within an interval $[0,\tau]$ in nonequilibrium steady-state processes were obtained:
$\begin{gathered}\Pr[S_{\mathrm{max}}(\tau)\geq s_{+}]\leq e^{-ps_{+}}\left<e^{pS_{\mathrm{tot}}(\tau)}\right>,\\ \Pr[S_{\mathrm{min}}(\tau)\leq-s_{-}]\leq e^{-ps_{-}}\left<e^{-pS_{\mathrm{tot}}(\tau)}\right>,\end{gathered}$ (14)
where the short-hand notation $S_{\mathrm{max}}(\tau)=\max_{t\leq\tau}S_{\mathrm{tot}}(t)$ and $S_{\mathrm{min}}(\tau)=\min_{t\leq\tau}S_{\mathrm{tot}}(t)$, is used, $s_{+},s_{-}>0$ are a given threshold, and $p\geq 1$ is a real number leading to a one-parameter family of bounds. These bounds place restrictions on the probability that the maximum and minimum excursions of the entropy production in the interval $[0,\tau]$ overcome the thresholds $s_{+}$ and $-s_{-}$, respectively. Following the original work [45], the probability bounds in Eq. (14) can be translated into optimal thresholds on the maximum and minimum entropy production that would not be surpassed with a given confidence probability during the interval. These bounds can also be used to bound the maxima and minima of the dissipated heat and the output work performed by steady-state heat engines and refrigerators.
In the case of molecular motors, one may observe the output mechanical work performed by the motor by consuming some chemical fuel,i.e. from the chemical work that comes from breaking molecular bounds, such as those of ATP. The procedure shown in Ref. [45] can be directly generalized to bounds for the output mechanical work performed by a molecular motor, as we show in Appendix C. As a result, we obtain optimal thresholds for the output mechanical work of the motor, which are guaranteed to not be overcome with a confidence probability $1-\alpha$, with $\alpha>0$ typically small:
@@
where $\eta$ is the macroscopic efficiency of the motor, $\eta=-W_{\mathrm{cycle}}^{\rm m}/W_{\mathrm{cycle}}^{\rm c}\leq 1$. This family of thresholds depends on the expectation values of $e^{\pm pS_{\mathrm{tot}}(\tau)}\right>$, making it cumbersome to perform the minimization on $p$. Nonetheless, in the long-term limit $t\to\infty$, the minimization can be performed and the lower threshold in Eq. (15) saturates to the following constant value (see Appendix C for details):
@@
which represents our third main relation for thermodynamic inference. In the last equation, assuming the chemical work per cycle $W_{\mathrm{cycle}}^{\rm c}$ is known, the mechanical work $W_{\mathrm{cycle}}^{\rm m}$ per cycle can be inferred from the current values overcoming the minimum threshold, $N_{-}^{*}=w_{-}/W_{\rm cycle}^{\rm m}$ in interval $[0,\tau]$ for a given value of $\alpha$, and environmental temperature $T$. Analogously, if $W_{\rm cycle}^{\rm m}$ is known but $W_{\rm cycle}^{\rm c}$ is not, one can estimate the latter from the minimum current.
@@
Figure 3: Three-states model for the F1-ATPase motor performing a cycle on the hydrolisis (anticlockwise direction) in three steps: Binding ATP ($x_{0}\rightarrow x_{1}$), hydrolising ATP into ADP + P, accompanied by a $120^{\circ}$ rotation against an external torque ($x_{1}\rightarrow x_{2}$), and releasing $ADP$ ($x_{2}\rightarrow x_{0}$). Only transitions that complete the cycle are explicit for the sake of readability, but the reverse transitions are also possible. The system and the reservoirs of ATP and ADP (colored squares) are all embedded in a dissolution at a particular temperature, $T$.
@@
[…]
@@
Table 1: Summary of the experimental values employed to characterize the rates of the system. Approximated concentrations taken from different experimental setups for these working parameters are $[\mathrm{ATP}]=[\mathrm{ADP}]=0.4\leavevmode\nobreak\ \mathrm{nM}$ and $[\mathrm{Pi}]=1\leavevmode\nobreak\ \mathrm{mM}$, and temperature $T=298\leavevmode\nobreak\ \rm{K}\leavevmode\nobreak\ (25^{\circ}\rm{C})$.
@@ III.1 Inference results from the exact IFT at stopping times
For F1-ATPase, we choose as a signaling transition the hydrolysis-synthesis transition ($x_{1}\leftrightarrow x_{2}$), which is the only transition that involves a macroscopically observable angular displacement of the motor $\Delta\theta=\pm 2\pi/3$ in one or the opposite directions. Following the stopping protocol introduced in Sec. II.1, we are interested in the net number of hydrolysis and synthesis transitions during a time interval, which can be obtained by recording the time evolution of the angular position of the motor. An illustration of the inference stopping protocol is shown in Fig. 1. After observing the first rotation in the hydrolysis direction at $t_{\rm onset}$, we track the angular position of the motor until a net number of hydrolysis $N_{+}$ or synthesis transitions $N_{-}$, that is, when the angular position of the motor reaches thresholds $(N_{+}+1)|\Delta\theta|$ or $-N_{-}|\Delta\theta|$, respectively. The heat that enters the system in a cycle in the hydrolysis direction is in this case:
@@
where the indexes $+$ and -- represent hydrolysis and synthesis, respectively.

### v2

@@ I Introduction
One of the most important molecular motors in which we focus in this work is the ATP synthase complex [1], responsible for the synthesis of ATP, the energy currency of life. It appears in practically all living organisms, from prokaryotic cells to eukaryotic cells. The complex is composed of two main parts: the F0 domain, embedded in the cellular membrane, and the F1 domain, also called F1-ATPase. The latter is a rotary motor capable of using the energy released by translocation of protons across the F0 domain to synthesize or hydrolize ATP molecules [16, 17]. The F1-ATPase motor has been widely studied since the end of the 90s, both experimentally [18, 19] and theoretically [20]. The thermodynamic operation of the motor has been accurately explored in the last decade, assessing some of the main features of the motor and suggesting a nearly 100% thermodynamic efficiency in its transduction of chemical energy to mechanical work [21, 22]. In this context, the so-called fluctuation theorem (FT) [23, 24] has been combined with single-molecule experiments to provide improved estimations of the rotatory torque in the motor [25], which constitutes one of the first applications of universal relations in stochastic thermodynamics for inference purposes [26, 27, 28]. This inference concept [29] has also been applied and extended to many other active biological processes, and to other universal non-equilibrium relations, such as the relation between dissipation and irreversibility [30, 31, 32, 33, 34, 35] , the thermodynamic uncertainty relation [36, 37, 38, 39], the thermodynamic speed limits [40, 41] or the infimum law of martingale theory [42, 43, 44].
@@ II Inference strategies using stopping times
In the following, we review relevant theoretical results on martingale theory for stochastic thermodynamics and introduce the two main inference strategies we propose in a generic configuration for discrete-state molecular motors performing cycles in its state space. For further details about the results of martingale theory in stochastic thermodynamic relevant to our strategies see Appendix A.
@@ II.1 IFT at stopping times
$\left<e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta(\mathcal{T})}\right>_{\mathcal{T}}=1,$ (1)
@@
$\left<e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta(\mathcal{T})}\right>_{\mathcal{T}}=1-\Gamma,$ (3)
@@ II.2 Counting cycles from observable transitions
Figure 1: Illustration of the stopping protocol introduced in the main text with sample trajectories from the F1-ATPase rotatory motor (see Sec. III). The current in the signalling transition $N(t)$ is monitored over time after the first transition is observed (after a time $t_{\rm onset}$), where we place the origin of times for the trajectories. Trajectories are stopped either when the current reaches the threshold value $N_{+}+1$ (blue trajectory) performing $N_{+}$ cycles in the forward direction, or $N_{-}$ (orange trajectory) corresponding to $N_{-}$ cycles in the backward direction after the starting point at $t_{\rm onset}$.
The stopping condition we propose consists on a double threshold first-passage time for the current $N(t)$, as illustrated in Fig. 1: the dynamics stops when the net number of signaling transitions performed by the motor reaches either the upper threshold $N_{+}+1$ or the lower threshold $N_{-}$, with positive numbers $N_{+}\neq N_{-}$ in general. We settle the starting point of the trajectories $X_{[0,\tau]}$ after the first signalling transition in the forward direction is detected, where we place the origin of time ($t_{\rm onset}$ in Fig. 1). We then let evolve the motor until the net number of jumps in the signalling transition reaches either the threshold $N_{+}+1$ or $N_{-}$, which corresponds to having trajectories with $N_{+}$ net transitions in the forward direction (i.e. from the starting point) or $N_{-}+1$ transitions in the backward direction, respectively. Moreover, here we will consider a limit time $\tau$ large enough to guarantee that the stochastic dynamics always verify the stopping condition before reaching the limit time, i.e. $\mathcal{T}<\tau$. This condition is ensured since the functional $e^{-S_{\mathrm{tot}}(\mathcal{T})-\delta_{\tau}(\mathcal{T})}$ remains bounded in the limit $\tau\rightarrow\infty$ [55], as guaranteed by the two-threshold first-crossing condition considered here.
@@
$\left<e^{-S_{\mathrm{tot}}(\mathcal{T})}\right>_{\mathcal{T}}=P_{-}e^{-S_{\mathrm{tot}}^{-}}+P_{+}e^{-S_{\mathrm{tot}}^{+}}=1-\Gamma,$ (6)
@@
$\begin{split}\Gamma&=\int_{0}^{\tau}d\mathcal{T}P(\mathcal{T})\sum_{X_{[0,\mathcal{T}]}\in AI}{\mathbb{P}}({\tilde{X}_{[\mathcal{T},0]}}|\mathcal{T})\frac{\pi_{x_{\mathcal{T}}}}{\rho_{x_{\mathcal{T}}}(\mathcal{T})}\\ &\simeq\int_{0}^{\infty}d\mathcal{T}P(\mathcal{T})\sum_{x(0)\neq x_{2}}\pi_{x(0)}=1-\pi_{x_{2}},\end{split}$ (7)
@@
On the other hand, when the boundary $N_{-}$ is reached, the entropy production becomes:
@@
where $\Delta E_{x_{2}x_{1}}=E_{x_{1}}-E_{x_{2}}$ is the energy difference between mesostates $x_{1}$ and $x_{2}$, $\Delta N_{x_{2}x_{1}}^{\alpha}$ is the number of particles $\alpha$ exchanged with the reservoir in the transition, and $W_{x_{2}x_{1}}^{\rm ext}$ is a contribution to the mechanical work performed by the system against an external force pointing in the $x_{2}\rightarrow x_{1}$ direction.
@@
$\left<e^{-S_{\mathrm{tot}}(\mathcal{T})}\right>_{\mathcal{T}}=$ $(1-P_{+})\pi_{x_{1}}e^{\beta\left[q_{x_{2}x_{1}}-N_{-}Q_{\mathrm{cycle}}\right]}$
@@
Remarkably, if the experimenter has access to the steady distribution values $\pi_{x_{1}}$ and $\pi_{x_{2}}$ (approximating $\Gamma\simeq 1-\pi_{x_{2}}$) and $q_{x_{1}x_{2}}$ is assumed to be known 111 In particular, in equilibrium, the local detail balance of each pair of transitions may be used to asses the energy (or free energy) changes between mesostates [29]. Moreover, $\pi_{x_{1}}$ and $\pi_{x_{2}}$ might be obtained in the long-time limit as the fraction of time the system spends in each state, which may require extra auxiliary monitorable transitions from $x_{1}$ and $x_{2}$., the heat per cycle $Q_{\mathrm{cycle}}$ can be inferred from Eq. (II.2), which may be used to obtain two main quantities: First, if the external forces acting on the system are known, one can infer the chemical work performed by the motor, and hence the corresponding differences in chemical potentials. Second, if all the chemical potentials are already known (as we will mainly consider in the following), one could use this information to measure external forces in the nanoscale.
@@ II.3 Inference from IFT with approximations
Nevertheless, in many situations of experimental interest only a few transitions among system states are observable. In that case, neither the stationary distribution of the states, $\pi_{x}$, nor the term $\Gamma$ would be accessible. To overcome this drawback, we propose a slightly different counting strategy and allow for some approximations. In this case, we assume to start the process directly from the stationary distribution $\pi_{x}$ in order to avoid absolute irreversibility. Therefore, now we start to count instances of the signalling transition without knowing exactly the initial state $x(0)$ (i.e. we do not wait any more for the first transition in the forward direction to happen to start counting). We then count the number of jumps in the signalling transition until the net flux $N(t)$ reaches one of the two thresholds $N_{+}$ or $N_{-}$. We remark that in this scenario, both $\delta(\mathcal{T})$ and $\Gamma$ become zero. Besides, we approximate the entropy production at the two thresholds by:
@@ II.4 Bounds to maximal excursions of entropy production
In Ref. [48], universal families of bounds for the cumulative distribution of entropy production maxima and minima (and thus for the survival probability) within an interval $0,\tau]$ in nonequilibrium steady-state processes were obtained:
[…]
where the short-hand notation $S_{\mathrm{max}}(\tau)=\max_{t\leq\tau}S_{\mathrm{tot}}(t)$ and $S_{\mathrm{min}}(\tau)=\min_{t\leq\tau}S_{\mathrm{tot}}(t)$, is used, $s_{+},s_{-}>0$ are a given threshold, and $p\geq 1$ is a real number leading to a one-parameter family of bounds. These bounds place restrictions on the probability that the maximum and minimum excursions of the entropy production in the interval $0,\tau]$ overcome the thresholds $s_{+}$ and $s_{-}$, respectively. Following the original work [48], the probability bounds in Eq. (14) can be translated into optimal thresholds on the maximum and minimum entropy production that would not be surpassed with a given confidence probability during the interval. These bounds can also be used to bound the maxima and minima of the dissipated heat and the output work performed by steady-state heat engines and refrigerators.
In the case of molecular motors, one may observe the output mechanical work performed by the motor by consuming some chemical fuel,i.e. from the chemical work that comes from breaking molecular bounds, such as those of ATP. The procedure shown in Ref. [48] can be directly generalized to bounds for the output mechanical work performed by a molecular motor, as we show in Appendix B. As a result, we obtain optimal thresholds for the output mechanical work of the motor, which are guaranteed to not be overcome with a confidence probability $1-\alpha$, with $\alpha>0$ typically small:
@@
where $\eta$ is the macroscopic efficiency of the motor, $\eta=-W_{\mathrm{cycle}}^{\rm m}/W_{\mathrm{cycle}}^{\rm c}\leq 1$. This family of thresholds depends on the expectation values of $\left<e^{\pm pS_{\mathrm{tot}}(\tau)}\right$>, making it cumbersome to perform the minimization on $p$. Nonetheless, in the long-term limit $t\to\infty$, the minimization can be performed and the lower threshold in Eq. (15) saturates to the following constant value (see Appendix B for details):
@@
which represents our third main relation for thermodynamic inference. In the last equation, assuming the chemical work per cycle $W_{\mathrm{cycle}}^{\rm c}$ is known, the mechanical work $W_{\mathrm{cycle}}^{\rm m}$ per cycle can be inferred from the current values overcoming the minimum threshold, $N_{-}^{*}=w_{-}/W_{\rm cycle}^{\rm m}$ in interval $0,\tau]$ for a given value of $\alpha$, and environmental temperature $T$. Analogously, if $W_{\rm cycle}^{\rm m}$ is known but $W_{\rm cycle}^{\rm c}$ is not, one can estimate the latter from the minimum current.
@@
Figure 3: Three-states model for the F1-ATPase motor performing a cycle on the hydrolisis (anticlockwise direction) in three steps: Binding ATP ($x_{0}\rightarrow x_{1}$), hydrolising ATP into ADP + P, accompanied by a $120^{\circ}$ rotation against an external torque ($x_{1}\rightarrow x_{2}$), and releasing ADPADP ($x_{2}\rightarrow x_{0}$). Only transitions that complete the cycle are explicit for the sake of readability, but the reverse transitions are also possible. The system and the reservoirs of ATP and ADP (colored squares) are all embedded in a dissolution at a particular temperature, $T$.
@@
The values for all parameters used in the simulations are given in Tab. 1. Their experimental validation is explained in Appendix C. We note that with these parameters the time needed to reach the stationary state is, on average, an order of magnitude shorter than the time needed to perform a single rotation in the hydrolysis direction (approximately $0.1\leavevmode\nobreak\ \text{s}$), as shown in Appendix D. This implies that the motor is always rotating in the stationary state, which guarantees the applicability of our approach.
@@
Table 1: Summary of the experimental values employed to characterize the rates of the system. Approximated concentrations taken from different experimental setups for these working parameters are $\mathrm{ATP}]=[\mathrm{ADP}]=0.4\leavevmode\nobreak\ \mathrm{nM}$ and $\mathrm{Pi}]=1\leavevmode\nobreak\ \mathrm{mM}$, and temperature $T=298\leavevmode\nobreak\ \rm{K}\leavevmode\nobreak\ (25^{\circ}\rm{C})$.
@@ III.1 Inference results from the exact IFT at stopping times
For F1-ATPase, we choose as a signaling transition the hydrolysis-synthesis transition ($x_{1}\leftrightarrow x_{2}$), which is the only transition that involves a macroscopically observable angular displacement of the motor $\Delta\theta=\pm 2\pi/3$ in one or the opposite directions. Following the stopping protocol introduced in Sec. II.1, we are interested in the net number of hydrolysis and synthesis transitions during a time interval, which can be obtained by recording the time evolution of the angular position of the motor. An illustration of the inference stopping protocol is shown in Fig. 1. After observing the first rotation in the hydrolysis direction at $t_{\rm onset}$, we track the angular position of the motor until a net number of hydrolysis $N_{+}$ or synthesis transitions $N_{-}$, that is, when the angular position of the motor reaches thresholds $(N_{+}+1)|\Delta\theta|$ or $N_{-}|\Delta\theta|$, respectively. The heat that enters the system in a cycle in the hydrolysis direction is in this case:
@@
where the indexes ++ and - represent hydrolysis and synthesis, respectively.